the sum of 2 consecutive reciprocals of 2 consecutive even integers is 9/40. what are the integers?
Answer by stanbon(75887)

the sum of 2 consecutive reciprocals of 2 consecutive even integers is 9/40. what are the integers?
----------------
Integers:
1st: 2x
2nd: 2x+2
----
Reciprocals:
1st: 1/(2x)
2nd: 1/(2x+2)
--------------------------
Equation:
1/(2x) + 1/(2x+2) = 9/40
---
Multiply thru by 40x(x+1) to get:
20(x+1) + 20x = 9x(x+1)
40x+20 = 9x^2+9x
----
9x^2-31x-20 = 0
Positive solution: x = 4
================================
1st: 2x = 8
2nd: 2x+2 = 10
===================
Cheers,
Stan H.
=============